# The anonymization problem in social networks

### v1

@@ 6.1 Experimental setup and data
We implement the algorithms introduced in Section 5 in our reusable C++ ANO-NET framework. [^1] https://github.com/RacheldeJong/ANONET For the preliminary experiments on graph models we used python and NetworkX [50].
We executed experiments on both graph models and real-world networks. In each experiment all edges are deleted. The recompute gap was chosen so that the last edges are deleted in the 100th step. Including the initial anonymity measurement this yields 101 data points for each algorithm. For the second and third variant, we set the target uniqueness to 95% of the nodes and the budget to 5% of the edges respectively. To account for the nondeterminism in both the algorithms and the generation of model networks, we report averages over five runs. For results showing the anonymization process, we additionally report the results ± one standard deviation.
@@ 6.2 Alteration operations and anonymity
Here we study the effectiveness of deleting, adding, and rewiring edges for the purpose of anonymization. Figure 1 shows results on how each operation affects the anonymity in graph models.
A first detail to note is that the number of unique nodes when no edges are altered, varies for the average degrees. Graphs with higher average degree have a higher initial uniqueness, which corresponds to findings in previous works [6, 7].
@@
The main conclusion from these results is that deletion is more effective than addition or rewiring, confirming what we theorized in Section 5.3. Logically, when all edges are deleted, this results in a graph without edges, making all nodes anonymous. On the contrary, edge addition makes the graph more dense, while for rewiring density remains constant. In many cases addition and rewiring result in a higher uniqueness, which is especially clear for the BA model, and average degree larger than 16 for both the ER and WS models. However, for the WS model, with average degree 16, rewiring performs slightly better than deletion when altering a fraction of 0.2 to 0.5 of the edges. This is likely due to the many nearly identical near-cliques in the generated networks.
While these preliminary results can not rule out that addition and rewiring can be valuable in specific cases, we choose to focus on edge deletion in the remainder of the paper.
Figure 2: Uniqueness when using different anonymity measures (color) and anonymization algorithms (linestyle) on real-world networks. For each network we show the fraction of deleted edges (horizontal axis) and the attained uniqueness (vertical axis).
@@ 6.3 Anonymization and measures
Figure 2 shows how the uniqueness changes when deleting edges using different measures for anonymity, for each of the seven anonymization algorithms. We depict a selection of five networks representative of the overall observed behaviour for the 17 networks listed in Table 2. This is, from left to right in Figure 2, a plateau with a steep decrease, linear decrease, plateau with linear decrease, sigmoid-like decrease and initial increase before decrease. For completeness, the results for the remaining 12 networks are included in Appendix B of the Supplementary materials.
@@
Overall, we see that the measure for anonymity used has a large influence on the results; both on the starting uniqueness with which the process starts and on how quickly the network is anonymized. The simplest measure, degree, has the lowest initial uniqueness and is the easiest to anonymize for. The measure with the furthest reach, vrq, has the highest initial uniqueness, which corresponds to the findings of [25], and is the most difficult to anonymize for.
@@
An example where different measures have the same initial uniqueness but differ in how quickly they are anonymized is the ‘Radoslaw emails’ network. For this network, the uniqueness starts around 0.78 for both count and $d$-$k$-anonymity, but decreases substantially faster for count. Making all nodes anonymous, however, requires a similar amount of edge deletions for both measures.
@@
In the ‘GRQC collab’ network, the results for some anonymization algorithms show an increase before decreasing. This is likely due to the collaboration network’s many cliques. When randomly deleting edges, these cliques are destroyed, which may initially result in higher uniqueness. If edges are targeted more specifically, e.g., using the u-aff-u heuristic, this increase does not occur.
@@
In the remainder of this paper, we focus on the measure count, as this measure is a realistic real-world scenario, has a substantial initial uniqueness compared to the other measures, yet is easier to anonymize for than for example the $d$-$k$-anonymity measure.
@@
6.4 Full and partial network anonymization
We now turn to the first and second variant of the anonymization problem, aiming to ensure anonymity for 100% and 95% of the nodes in the network. The top of Figure 3 shows which fraction of edges can be preserved when deleting edges until the desired level of anonymity is reached. This fraction equals $\frac{|E^{\prime}|}{|E|}$ where $E$ is the set of edges in the original network, and $E^{\prime}$ in the anonymized network. Here, a higher fraction indicates better results, as preserving more edges overall implies a higher utility of the resulting anonymized network. For the partial variant, both ‘US Power grid’ and ‘Euroroads’ retain 100% of their edges as initially less than 5% of their nodes is unique (so, 95% is already anonymous).
The results show that for some networks, such as ‘Radoslaw emails’ and ‘FB Reed98’ only a small fraction of edges can be preserved, which is also clear from results in Section 6.3. For other networks, larger fractions can be preserved. In these cases, the uniqueness based algorithms, specifically aff-u and u-aff-u, clearly outperform the other algorithms. When dividing the result for the best heuristic by the result of es, we obtain the baseline improvement ratio reported in Table 2. For the full and partial variant, on average 17 and 3.0 times more edges can be preserved.
Moving to the other algorithms, the results show that the structure based algorithms perform similar. For the uniqueness based algorithms, aff-u and u-aff-u perform similar and always achieve better results than the es baseline. However, unique performs worse than the other uniqueness based algorithms.
Turning to the results for anonymizing 95% of the nodes, we see that for some networks a much higher fraction of edges can be preserved. Examples are ‘Moreno health’ and ’Pajek Erdos’. This indicates that these networks contain nodes that are difficult to anonymize and require much attention. When this is the case, one could consider to let more nodes remain unique in order to preserve more edges and therewith achieve an overall higher utility of the resulting anonymized network.
@@ 6.5 Budgeted anonymization
We now move to results for the third variant, being budgeted anonymization. As the networks each start at different uniqueness values, as can be seen in Table 2, the values on the vertical axis on the bottom of of Figure 3 show the fraction of unique nodes that is anonymized, i.e., $1-\frac{uniqueness(G^{\prime})}{uniqueness(G)}$, when deleting at most 5% of the edges. A higher value thus indicates that the algorithm is more effective in anonymization within the given budget.
@@
For some of the networks, the anonymization algorithms are quite effective. As can be derived from Table 2, this overall holds for networks with lower initial uniqueness and higher average distance. Experiments to confirm this are included in Appendix C of the Supplementary materials. However, some of the networks, such as ‘Radoslaw emails’ and ‘GRQC collab’, require more than 5% deletion to obtain a substantial decrease in uniqueness. This is also visible in Figure 2 showing a plateau for ‘Radoslaw emails’ and larger initial increase in uniqueness for ‘GRQC collab’.
@@
Overall, we see a big performance difference both between networks and between anonymization algorithms. For some of the networks, such as ‘Moreno health’, ‘Euroroads’, ‘Copnet sms’ or ‘US power grid’, the two best uniqueness based anonymization algorithms, aff-u and u-aff-u, manage to anonymize many more nodes compared to the other algorithms. For all networks, these two uniqueness based algorithms either perform similar or outperform the other algorithms. For the ‘Euroroads’ network, this even results in all nodes being anonymized within the given budget. These algorithms also showed to be most effective for the full and partial variants. On average, 4.4 times more nodes are anonymized compared to the baseline es algorithm. The algorithms especially perform well for networks with a lower initial uniqueness, lower average degree and higher average distance, which is further explained in Appendix C of the Supplementary materials.
Figure 4: Runtimes for anonymization algorithms (color) with $b=|E|$ and recompute gap $|E|/100$.
6.6 Runtimes
Figure 4 shows the observed runtime of each algorithm on each network by our ANO-NET framework. The runtime consists of both computing probabilities to choose edges and bookkeeping operations, including updating the partition and deleting the chosen edges from the network (line 1 and 1 in Algorithm 1).
The results show that the runtimes vary considerably between networks, which are sorted based on the number of nodes. The first two networks, ‘Radoslaw emails’ and ‘Primary school’, which contain the smallest number of nodes, require more time than the ones following in node size, ‘Moreno innov.’ and ‘Gene fusion’. This is likely due to their high uniqueness throughout the process which makes anonymization more computationally expensive.
@@ 6.6 Runtimes
Moving to the different algorithms, we first note that the es baseline spends a minimum amount of time on edge selection and hence most time is spent on bookkeeping. Overall, there is no large difference in the runtimes achieved by the algorithms. When comparing the runtimes more closely, uniqueness based algorithms are often slightly more time consuming, as computation of the probability assigned to edges is more expensive. However, for some networks, e.g., ‘Air traffic control’ and ‘Network science’, u-aff-u is slightly faster than the other algorithms, including es. This can likely be explained by the fact that u-aff-u is more effective in anonymization, which results in fewer unique nodes early in the anonymization process. This, in turn, results in lower runtimes for the bookkeeping operations and the edge probability computation.
@@ 7 Conclusion
In this paper, we introduced the anonymization problem in networks in the form of three variants being 1) full, 2) partial and 3) budgeted anonymization. We introduced a generic framework coined ANO-NET which allows one to anonymize networks with various settings, including measures for anonymity, different anonymization algorithms and relevant parameters.
Of the anonymization operations that can be applied, results on graph models showed that edge deletion is most effective. We proposed and experimentally evaluated seven measure-agnostic anonymization algorithms, categorized into edge sampling, structure based and uniqueness based approaches. Experiments to assess the effect of different anonymity measures on the anonymization process highlighted how networks are more difficult to anonymize using a more strict measure. The choice in measure affects both the initial uniqueness and how quickly uniqueness decreases.
The effectiveness of the algorithms differed noticeably per network. In particular, networks with lower initial uniqueness and higher average path length, showed to be easier to anonymize. From the anonymization algorithms two of the uniqueness based algorithms which account for the number of unique nodes affected, consistently showed to be most effective for anonymization. Compared to the baseline, the best heuristic algorithm anonymized on average 4.4 times more nodes for the budgeted variant, and preserved 17 and 3.0 times more edges for the full and partial variant. While overall comparable, runtimes for the uniqueness based algorithms appear affected by the initial uniqueness and speed of uniqueness decrease obtained.
The aim of this paper was to present a general version of the anonymization problem and its variants, study various properties of the problem, and evaluate seven relatively simple heuristic algorithms. An obvious next step is to design (metaheuristic) anonymization algorithms, or explore the use of machine learning. Additionally, it would be interesting theoretically to determine the complexity of the problem when using measures such as count or vrq. Lastly, a step forward would be to dynamically account for how data utility decreases during the anonymization process.
Acknowledgments
This research was made possible by the Platform Digital Infrastructure SSH (http://www.pdi-ssh.nl). We would also like to thank the POPNET team (https://www.popnet.io) and the Leiden CNS group (https://www.computationalnetworkscience.org) for various helpful suggestions and discussions.

### v2

@@ 6.1 Experimental setup and data
We implement the algorithms introduced in Section 5 in our reusable C++ ANO-NET framework. [^2] https://github.com/RacheldeJong/ANONET For the preliminary experiments on graph models we used python and NetworkX [61], and for utility analysis we used igraph [62].
We executed experiments on graph models and real-world networks. In each experiment, all edges are sequentially deleted unless the process exceeds the runtime limit of 30 minutes. In this case, the process continues only until the required number of edges is deleted for the budgeted variant. For the four networks that this pertains to, only results on observed runtime are reported, with the goal of assessing scalability of the anonymization algorithms. For the partial and budgeted variant, we set the target anonymity to 95% of the nodes and the budget to 5% of the edges respectively. The budget is chosen based on previous work indicating that randomly deleting 5% of the edges preserves most network properties [41, 54].
The recompute gap was chosen so that the last edges are deleted in the 100th step. Including the initial anonymity measurement this yields 101 data points for each run. To account for the nondeterminism in both the anonymization algorithms and the generation of model networks, we report averages over five runs. For results showing the anonymization process and utility analysis, we additionally report the standard deviation.
For results on utility, we generate for each run the anonymized graph for all 101 data points by deleting the selected edges, and for each compute the utility metrics summarized in Section 5.4. For results on community detection, we run the Leiden algorithm [5] 10 times for each network, followed by a consensus clustering step [51].
@@ 6.2 Alteration operations and anonymity
Here we study the effectiveness of deleting, adding, and rewiring edges for the purpose of anonymization. Figure 1 shows results on how repeatedly applying each operation at random affects the anonymity in graph models.
A first detail to note is that the number of unique nodes when no edges are altered, varies with the average degree. Graphs with higher average degree have a higher initial uniqueness, which corresponds to findings in previous works [7, 8].
The main conclusion from these results is that random deletion is more effective than random addition or rewiring, confirming what we theorized in Section 5.3. Logically, when edges are deleted, this results in a graph with lower density with less variety in neighborhoods, which could in turn result in more anonymous nodes. On the contrary, edge addition makes the graph more dense, resulting in larger neighborhoods, while for rewiring density remains constant. For the graph models, random addition and rewiring can even result in a higher uniqueness, which is especially clear for the BA model, and for the ER and WS models when the average degree is larger than 16. However, for the WS model, with average degree 16, rewiring performs slightly better than deletion when altering a fraction of 0.2 to 0.5 of the edges. This is likely due to the many nearly identical near-cliques in the generated networks.
While these preliminary results can not rule out that edge addition and rewiring can possibly be more effective in specific cases, and could be preferred for certain applications, we choose to focus on edge deletion in the remainder of the paper.
@@ 6.3 Anonymization and measures
Figure 2 shows how the uniqueness changes when deleting edges using different measures for anonymity. For readability, the results are shown for a subset of the anonymization algorithms, with one algorithm for each category in Section 5.3. We depict a selection of five networks representative of the overall observed behavior for the networks listed in Table 2. This is, from left to right in Figure 2, 1) a plateau with a steep decrease, 2) linear decrease, 3) plateau with linear decrease, 4) sigmoid-like decrease and 5) initial increase before decrease. The results for the remaining networks are included in Appendix B.
@@
Overall, we see that the measure for anonymity used has a large influence on the results; both on the starting uniqueness with which the process starts and how fast the uniqueness decreases. The simplest measure, degree, has the lowest initial uniqueness and is the easiest to anonymize for. The measure with the furthest reach, vrq, has the highest initial uniqueness, which corresponds to the findings of [26], and is the most difficult to anonymize for, even more difficult than $d$-$k$-anonymity which accounts for exact structure of the neighborhood of nodes.
@@
An example where different measures have the same initial uniqueness but differ in how quickly they are anonymized is the “Radoslaw emails” network (left). For this network, the uniqueness starts around 0.78 for both count and $d$-$k$-anonymity, but decreases substantially faster for count. However, making all nodes anonymous, i.e., full anonymization, requires a similar amount of edge deletions for both measures.
@@
In the “GRQC collab” network (right), the results for the es algorithm shows an increase in uniqueness before decreasing. This is likely due to the collaboration network’s many cliques. When randomly deleting edges, these cliques are destroyed, which may make certain nodes unique. If edges are targeted more specifically, e.g., using the ua heuristic, this initial increase does not occur.
@@
In the remainder of this paper, we focus on the measure count, as this measure is a realistic real-world scenario, has a substantial initial uniqueness compared to the other measures, yet is easier to anonymize for than, for example, the $d$-$k$-anonymity measure. Using the count measure, nodes are essentially equivalent if they have the same number of neighbors, and are part of the same number of triangles.
Figure 2: Uniqueness when using four different anonymity measures (color) and three anonymization algorithms (linestyle) on real-world networks. For each network we show the fraction of deleted edges (horizontal axis) and the attained uniqueness (vertical axis).
6.4 Full and partial network anonymization
In the following, we investigate how effective the proposed algorithms from Section 5.3 are in ensuring anonymity for 100% (full anonymization) and 95% (partial anonymization) of the nodes in the network. The top of Figure 3 shows the fraction of edges preserved $\frac{|E^{\prime}|}{|E|}$ when edges are deleted until the desired level of anonymity is reached. Here, a higher fraction indicates better results, as, in general, preserving more edges overall implies a higher utility of the resulting anonymized network. We look at utility in detail in Sections 6.6 and 6.7. For the partial variant eight of the networks, including “US Power grid” and “Euroroads” retain 100% of their edges as initially less than 5% of their nodes is unique.
The results show that for some networks, such as “Radoslaw emails” and “FB Simmons81” only a small fraction of edges can be preserved. For other networks, for which more edges can be preserved, the uniqueness based algorithm ua outperforms the other algorithms. When dividing the result for the best heuristic by the result of es, we obtain the improvement ratio reported in the three rightmost columns of Table 2. For the full and partial variant, on average 13.9 and 1.8 times more edges can be preserved.
@@ 6.5 Budgeted anonymization
We now move to results for budgeted anonymization. To account for the diverse values in starting uniqueness, as indicated in Table 2, the values on the vertical axis on the bottom of Figure 3 show the fraction of unique nodes that is anonymized, i.e., $1-\frac{U(G^{\prime})}{U(G)}$, when deleting at most 5% of the edges. A higher value thus indicates that the algorithm is more effective in anonymization within the given budget.
@@
For some of the networks, the ua algorithm is quite effective. As can be derived from Table 2, this overall holds for networks with lower initial uniqueness and higher average distance. Experiments to confirm this are included in Appendix C. However, some of the networks, such as “Radoslaw emails” and “GRQC collab”, require more than 5% deletion to obtain a substantial decrease in uniqueness. This can also be derived from Figure 2 showing a plateau for “Radoslaw emails” and an initial increase in uniqueness for “GRQC collab”.
@@
Overall, we see a big performance difference both between networks and anonymization algorithms. For some of the networks, such as “Moreno health”, “Euroroads”, “Copnet sms” or “US power grid”, the uniqueness based anonymization algorithm, ua, manages to anonymize many more nodes compared to the other algorithms. For all networks, this algorithm either performs similar or outperforms the other algorithms. For the “Euroroads” network, ua even succeeds to anonymize all nodes within the given budget. On average, 4.8 times more nodes are anonymized compared to the baseline es algorithm. The algorithms especially perform well for networks with a lower initial uniqueness, lower average degree and higher average distance, which is further explained in Appendix C.
6.6 Utility
In this section we investigate the utility of the anonymized networks in terms of how well various network properties and performance on downstream tasks are preserved for the different variants of the anonymization problem. Figure 4 summarizes this for the five anonymization algorithms. A property is considered preserved if the value after anonymization $\pm$ the standard deviation differs less than 5% from the original value. As discussed in Section 6.1, we exclude the four largest networks. Additionally, for partial anonymization, we exclude the eight networks which are already partially anonymous, i.e., networks with an initial uniqueness less than 5%, resulting in 11 networks for this variant.
When comparing results for the three variants of the problem, we observe that for full and partial anonymization properties are usually not preserved, except for the largest component size which the ua algorithm manages to preserve this for six networks for the full variant. This difference compared to other algorithms is likely due to the fact that this algorithm requires fewer edge deletions. Moving to the budgeted variant, we see that more often properties are preserved, and there is a small difference in results between the algorithms. Hence, focusing on the budgeted variant seems like a promising way to further investigate and compare anonymization algorithms.
Focusing on the different properties, we see that the largest component size is typically preserved best: for budgeted anonymization, ua even manages to preserve this property for all networks. This is also shown in Appendix D, which shows for each utility metric the difference to the initial value changes when deleting more edges. The average shortest path length and most central nodes according to betweenness centrality are the second and third best preserved properties. The clustering coefficient is not preserved for most networks and out of the compared algorithms unique performs best. This could be due to the fact that this algorithm does not target edges that affect many nodes, which can potentially destroy many triangles. The community structure is not preserved by any algorithms for most networks. This property seems especially challenging to preserve, as deleting the first percentage of edges already results in a large difference, despite using the consensus clustering step to obtain more stable communities.
Overall, ua best preserves the properties for all variants. Compared to the other algorithms in the budgeted variant, it best preserves most properties, and performs similar in preserving the clustering coefficient and average shortest path length.
Figure 4: For each of the three variants of the anonymization problem (columns) the color and number indicate the fraction of networks for which the network property or performance on a downstream task (vertical axis) is preserved, i.e., the change $\pm$ one standard deviation is less than 5% of the original value, using the five anonymization algorithms (horizontal axis).
6.7 Anonymity vs. utility
After comparing algorithms in terms of achieved anonymity and utility, we now investigate the trade-off between the two. Figure 5 shows for two network datasets the Pareto optimal solutions in terms of uniqueness and performance on downstream tasks found by the algorithms. Hence, the figure shows only solutions such that there are no other solutions found by the same algorithm with both a lower uniqueness and lower difference in performance on the downstream task. For readability, we included two networks that show the overall behavior found in the full set of networks for which the results can be found in Appendix E.
When comparing the algorithms, we see that ua always gives either the best trade-off between utility and anonymity, or one similar to the best algorithm. Moreover, on the downstream tasks, the ua algorithm always gives strictly better values than es, which often gives the worst trade-off of all algorithms. For other networks, such as “FB Simmons83”, the values found by the algorithms are more similar and more Pareto optimal solutions are found, indicating many small improvements are made in uniqueness during the anonymization process.
For the different metrics, we observe two types of behavior. First, for the largest connected component size, no large difference is required for the best algorithms unless one wants to achieve small uniqueness values. For this task, many edge deletions can be made until a larger change occurs, which is also shown by the results in Appendix D. Especially if preserving the giant component is of high importance, it would be worth considering the partial variant of the problem.
Moving to community and betweenness centrality, we first observe that a large difference is lost in order to decrease uniqueness. This result can also be seen in Appendix D, where deleting the first percentage of edges results in a large difference in both the most central nodes and communities. After this decrease, the difference increases for both metrics, but grows less quickly for the community aspect.
Figure 5: Pareto optimal solutions in terms of uniqueness (horizontal axis) and performance on downstream tasks (vertical axis). Each dot represents a solution on the Pareto front found by an anonymization algorithm (color). Filled dots represent solutions for which less than 5% is deleted.
6.8 Runtime
Figure 6 shows the observed runtime of each algorithm, on each of the three variants, for each network. The runtime consists of computing probabilities to choose edges, for which the computational complexity varies for the anonymization algorithms introduced in Section 5.3, deleting the chosen edges from the network, which can be done in linear time, and accordingly determining the new anonymity of nodes by updating the equivalence partition (lines 6 and 7 in Algorithm 1). For this last operation, which is a computationally expensive part of the anonymization process, the runtime is affected by the number of affected nodes and the density of the network.
Comparing the different problem variants, the lowest runtimes are usually observed for the the budgeted variant, except for networks that do not require edge deletions for the partial variant if their initial uniqueness is less than 5%. Runtimes for the partial and full variant are often very similar. This small difference is likely due to the fact that later iterations require less runtime, which is further supported by results on the runtime per iteration in Appendix F.
Overall the runtime increases for networks with more nodes. At the same time, the first four networks show that a high uniqueness and density increase runtimes. The latter makes both measuring anonymity and computing the number of affected nodes more computationally expensive. For the four largest networks, which were included to assess the scalability of the algorithms, we observed runtimes of over 30 minutes for the budgeted variant for all algorithms, including es which spends a minimal amount of time on selecting edges.
@@ 6.8 Runtime
Moving to the different algorithms, we first note that for many of the networks, there is no large difference in the runtimes required by the algorithms, implying that even on large networks the runtime can be high due to the bookkeeping operations. As shown in Appendix F, the time required for bookkeeping operations, including updating the equivalence classes, is usually much larger than the time used for the anonymization algorithms. Exceptions are networks such as “FB Simmons” for which the unique and ua spend the largest amount of time on the anonymization algorithm itself. However, for some networks, such as “Air traffic control” and “Network science”, the ua algorithm is slightly faster than the other algorithms, including es. This can be explained by the fact that ua is more effective in anonymization, which results in fewer required iterations and a lower runtime.
Figure 6: Runtimes for the five algorithms (color) with recompute gap $|E|/100$. Runtimes are shown for the three variants, full (bar), partial (circle) and budgeted (triangle). For the four largest networks, only runtimes of the budgeted variant are included.
@@ 7 Conclusion
In this paper, we introduced the anonymization problem in social networks in the form of three variants being 1) full, 2) partial and 3) budgeted anonymization. We introduced a generic algorithmic framework coined ANO-NET which allows one to anonymize networks, implementing the three variants, four measures for anonymity, the proposed anonymization algorithms and relevant parameters.
Of the edge alteration operations that can be applied, results on graph models showed that random edge deletion is most effective. Experiments to assess the effect of different anonymity measures on the anonymization process highlighted how networks are more difficult to anonymize using a more strict measure. The choice in measure affects both the initial uniqueness and how quickly uniqueness decreases. We proposed and experimentally evaluated five (of which four new) measure-agnostic anonymization algorithms. The effectiveness of the algorithms differed noticeably per network. In particular, networks with lower initial uniqueness and higher average path length showed to be easier to anonymize. From the anonymization algorithms one of the uniqueness based algorithms, ua, which accounts for the number of unique nodes affected, consistently showed to be most effective for anonymization. Compared to the baseline, this heuristic algorithm anonymized on average 4.8 times more nodes for the budgeted variant, and preserved 13.9 and 1.8 times more edges for the full and partial variant. Additionally, this algorithm managed to preserve three of the utility metrics better in budgeted anonymization. Furthermore, the ua algorithm appears to result in the best trade-off between anonymity and utility. While overall comparable, runtimes for the uniqueness based algorithms appear affected by the initial uniqueness and network density.
The aim of this paper was to lay the ground work for studying a general version of the anonymization problem and its variants, to study various properties of the problem, and evaluate five relatively simple heuristic algorithms both in terms of improving anonymity, and preserving data utility. An obvious next step is to design metaheuristic anonymization algorithms, or to explore the use of machine learning. One can also consider the use of the addition or rewiring operation, possibly in combination with deletion. To account for more attacker scenarios, it would be interesting to research which network properties, which may not be captured by the anonymity measures, could be used for de-anonymization, and design extensions accordingly. These extensions could both be based on structure, or account for additional network properties such as node or edge labels, or even temporality. Lastly, a step forward would be to dynamically account for how data utility decreases during the anonymization process and decide to continue or stop the anonymization process on-the-fly.